OpenStudy (anonymous):

Express the rational expression in lowest terms. 3x-3y-bx+by/3x-3y+bx-by (The answer is 3-b/3+b I just need to see how the problem is done, thank you!)

OpenStudy (phi):

\[ \frac{3x-3y-bx+by}{3x-3y+bx-by }\] first do the top it looks like we can factor 3 from the first 2 terms 3x-3y = 3(x-y) also we can factor out a "b" from the last two terms: -bx+by= b(-x+y) we have (x-y) in the first term and (-x+y) in the 2nd but if we factor out a -1, we have -(x-y) so the top is 3(x-y) - b(x-y) next, treat (x-y) as one "thing" and factor it out we get (3-b)(x-y) now do something similar for the bottom